OpenStudy (faiqraees):

P(A) = 0.35 P(B) = 0.4 Probability of neither A nor B if a) A and B are mutually exclusive events b) A and B are independent events @ganeshie8 @vishweshshrimali5 @ParthKohli And how can a flipping of coin be dependent event?( because mutually exclusive = dependent event)

OpenStudy (anonymous):

Mutually Exclusive: \[A\cap B = \emptyset\] Independent: \[P(A|B)=P(A),P(B|A)=P(B)\] so... \[\frac{P(A\cap B)}{P(B)} = P(A|B)=P(A)\implies P(A\cap B) = P(A)\cdot P(?)\]

OpenStudy (faiqraees):

@Redcan I know all this. I just want to know how to solve the question

OpenStudy (faiqraees):

@nincompoop

OpenStudy (faiqraees):

@inkyvoyd

OpenStudy (bobo-i-bo):

Properties and axioms of Probability. If E and F are two events: \[1. P( \emptyset ) = 0\] \[2. P(E^c)=1-P(E)\] \[3. P(E \cup F) = P(E) + P(F) - P(E \cap F) \] These are enough for you to find your answers.

OpenStudy (superteenblue):

hello

OpenStudy (faiqraees):

Are the answers 1 and 0.86?

OpenStudy (bobo-i-bo):

Nope :P

OpenStudy (faiqraees):

Which is wrong?

OpenStudy (faiqraees):

@Bobo-i-bo

OpenStudy (bobo-i-bo):

Both

OpenStudy (faiqraees):

Really? Then can you show me how to solve it?

OpenStudy (faiqraees):

@Bobo-i-bo

OpenStudy (bobo-i-bo):

Soo... can I see your working out and attempt? That would be a good starting point ^_^

OpenStudy (faiqraees):

For a) A∩B = 0 A'UB' = (A∩B)' A'UB' = 1-A∩B A'UB' = 1-0 A'UB' = 1

OpenStudy (faiqraees):

@Bobo-i-bo

OpenStudy (bobo-i-bo):

So the error is, you attempted to find \[P(A^c \cup B^c) \] But you should be trying to find \[P((A \cup B)^c)\]

OpenStudy (faiqraees):

Wait a minute what is the Set notation of neither A nor B?

OpenStudy (bobo-i-bo):

My notation: \[A^c\] is the same as your notation of \[A' \]

OpenStudy (faiqraees):

I know that

OpenStudy (bobo-i-bo):

Try drawing venn diagrams for \[A^c \cup B^c \] and \[(A \cup B)^c\]

OpenStudy (bobo-i-bo):

That will make the answer to "Wait a minute what is the Set notation of neither A nor B?" evident

OpenStudy (faiqraees):

so 0.25?

OpenStudy (faiqraees):

@Bobo-i-bo

OpenStudy (bobo-i-bo):

yup

OpenStudy (bobo-i-bo):

:)